# thanks: current affiliation: IBM Quantum

### v1

@@ III.2 Logical error suppression and detection by LDD
For all our experimental results, ‘No DD’ refers to $[[4,2,2]]$ encoding without any physical or logical DD. In all other experiments, we use physical XY4 to pad all the encoding and unencoding circuits gaps, and combine them with various flavors of LDD. This choice allows us to clearly assess the improvements introduced by DD, compared to relying solely on the code’s error detection capabilities.
III.2.1 Logical $Z$ errors
@@
We start by gradually increasing the time delay $\tau$ (up to 55 $\mu$s) between the encoding and unencoding without DD. This situation is relevant in the context of QEC experiments. For example, one could prepare an encoded qubit and then leave it idle while other logical operations are applied to other encoded qubits [19].
@@
We perform the $\overline{Z}_{i}$ error detection using the logical Bell states |$\ket{\Psi_{+}}$ and |$\ket{\Phi_{+}}$ as discussed in Section III, and display the results in Fig. 3. As can be seen in Fig. 3(a) and (c), without LDD the free evolution fidelity (denoted as No DD) decays rapidly and exhibits $ZZ$ crosstalk-induced oscillations. Fig. 3(b) and (d) show that logical $Z$ errors accumulate over time. We next use two types of LDD sequences to suppress this effect: Logical XX (LXX) and Logical XY4 (LXY4). Physical-level schematics of these LDD sequences are shown in the boxed four-pulse sequences of Fig. 1(d) and (e), respectively (disregarding the tilde notation). Their logical-level counterparts are shown at the bottom of Fig. 1. We generate both LXX and LXY4 using the native logical operations of the $[[4,2,2]]$ code. For example, in the upper part of Fig. 1(d), reading the first column from top to bottom yields $XIXI=\overline{XX}$, which is the $\bar{X}\bar{X}$ column at the bottom of Fig. 1(d). Reading the pulse sequences from left to right, the staggering (appearance of delays as indicated by the identity operations) is deliberately introduced to suppress crosstalk at the physical level [65]; see Appendix B for more details.
@@
The corresponding results are shown as dashed lines in Fig. 3. At short times ($\lesssim 15\mu$s), these results are better than those without DD because some errors (including $ZZ$ crosstalk) are suppressed by both LXX and LXY4. However, at longer times, the benefit is lost and, moreover, small oscillations appear that indicate the presence of coherent errors  [99]. To overcome this, we create robust versions of LXY4 and LXX by ensuring that all physical qubits in the code undergo physical DD sequences robust to pulse errors. Specifically, we use the universally robust (UR) sequence family [62], and ensure that each physical qubit undergoes a UR4 sequence, i.e., $X\tilde{X}\tilde{X}X$ or $Y\tilde{Y}\tilde{Y}Y$, where a tilde denotes an $X$ or $Y$ rotation by $-\pi$ instead of $\pi$. These robust versions, which we call RLXX and RLXY4, are the full sequences shown in Fig. 1(d) and (e). The performance of these sequences is shown by the solid lines in Fig. 3 and exhibits a significant improvement. Notably, the logical Bell state fidelities decay more slowly and without oscillations (panels (a) and (c)), and the logical $Z$ errors are strongly suppressed (panels (b) and (d)).
@@
In addition, we apply the physical staggered XY4 (SXY4) sequence. This corresponds to applying a single XY4 sequence to each physical qubit but in a staggered manner to reduce crosstalk [65]. This sequence, shown in Fig. 1(f), also corresponds to an LDD sequence for the $[[4,2,2]]$ code (bottom of panel (f)). However, it is not a universal logical sequence, but rather a purely logical-$Z$-error suppressing sequence. Interestingly, Fig. 3 shows that SXY4 performs on par with RLXX and RLXY4. This finding signals that the dominant logical errors are of $Z$-type. We confirm this in the following.
@@ III.2.2 Logical $X$ and $Y$ errors
III.2.3 LDD with postselection
Having implemented LDD, we can further improve the results by using the error detection capability of the $[[4,2,2]]$ code, which allows us to perform postselection. We do so by discarding any measurement outcome outside the logical basis (which would be the result of a physical error); i.e., we only keep measurement results corresponding to the bitstrings $\{0000,0110,1001,1111\}$.
@@ III.2.3 LDD with postselection
III.2.4 Beyond breakeven and the state-of-the-art
Fig. 5 also includes results for physical (unencoded) Bell states. Here we show only the best Bell pair among all pairs we tested, both without DD and with staggered (crosstalk-robust [65]) XY4. The former (‘Best Physical No DD’) exhibits strong crosstalk-induced oscillations with an overall fidelity comparable to that of the mean logical encoded Bell pairs case without DD (No DD+PS). Adding SXY4 significantly improves the fidelity and outperforms even the mean fidelity of logical Bell pairs with LDD (RLXY4). This shows that LDD by itself is not better than working with physical qubits and a crosstalk-robust DD sequence. However, physical Bell pairs with SXY4 are significantly worse than RLXY4+PS, i.e., the case of LDD with postselection on the results of the $[[4,2,2]]$ code. This constitutes clear evidence of beyond-breakeven performance for our QEC-LDD strategy.
@@ III.2.4 Beyond breakeven and the state-of-the-art
Figure 6: The percentage of data discarded per circuit as a function of the delay time between encoding and unencoding (dataset 2). While LDD suppresses logical errors, it also reduces certain physical errors, as evidenced by the lower data discard rate when using LDD (i.e., fewer physical errors are detected).
III.3 Physical error suppression and detection by LDD
The LDD sequences suppress not only logical errors but physical errors as well. To see this, consider, e.g., the LXX sequence [Fig. 1(d)]. As described in Section III.2.1, reading the pulse sequence vertically, each time step of this sequence operates in the logical subspace as $\overline{XX}$. Simultaneously, reading the sequence horizontally over a complete round of LDD, all four physical qubits undergo the physical XX sequence (i.e., $X-\tau-X-\tau$), which suppresses the set of physical errors $\{Y_{i},Z_{i}\}_{i=1}^{4}$. Similarly, for the RLXY4 sequence [Fig. 1(e)], each physical qubit undergoes UR4. which robustly [62] suppresses the sets of physical errors $\{Y_{i},Z_{i}\}_{i=1,3}$ (due to the $X\tilde{X}\tilde{X}X$ sequence) and $\{X_{i},Z_{i}\}_{i=2,4}$ (due to the $Y\tilde{Y}\tilde{Y}Y$ sequence).
One way to gauge the impact of this suppression of physical errors is shown in Fig. 6, which displays the percentage of discarded data per circuit as a result of postselection over $55\mu$s (out of $4000$ shots), comparing No DD to RLXY4 for the two logical Bell states |$\ket{\Phi_{+}}$ and |$\ket{\Psi_{+}}$. The results are obtained through bootstrapping across the ten sets of four qubits used in these experiments. Significantly less data are discarded with RLXY4 than without DD. Since discarded data correspond to the detection of physical errors, this means that LDD not only enhances the fidelity of logical Bell states by suppressing logical errors, but also reduces the occurrence of physical errors.
We can go further and use Algorithmic Error Tomography [54] to identify specific physical error types. For example, the $0100$ and $0010$ bitstrings correspond, respectively, to physical $Z_{2},Z_{4}$ and $Z_{1},Z_{3},$ errors in the |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ logical Bell state experiment. In Fig. 7 we show the corresponding relative bitstring counts, which are measures of detecting either of these errors. Additionally, we show the $0000$ and $0110$ relative bitstring counts, which correspond to the logical fidelity and the logical $Z$ error, respectively. Without LDD (dashed), we observe that the logical fidelity is low and both logical and physical errors increase and oscillate. With DD (solid), the RLXY4 sequence strongly suppresses both types of errors.
These results demonstrate that, as claimed, LDD sequences suppress both physical and logical errors.
Figure 7: The relative count of the $\{0000,0110,0100,0010\}$ bitstrings in the |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ experiments (dataset 2). The two even-weight bitstrings correspond to the logical fidelity ($0000$) and logical $Z$ error ($0110$), and the two odd-weight bitstrings correspond to the detection of physical $\{Z_{i}\}_{i=1}^{4}$ errors. Dashed (solid) lines correspond to no DD (RLXY4). The LDD sequence, in addition to suppressing the logical errors and improving the logical fidelity, also strongly suppresses the occurrence of these physical errors.
IV Discussion
The operation of QEC codes is adversely affected by the occurrence of logical errors that the code cannot detect or correct. Here, we have shown how to combine QEC with dynamical decoupling implemented in terms of the logical operators of the code, resulting in a hybrid QEC-LDD strategy that is significantly more effective than either QEC or LDD alone.
We designed our LDD sequences to simultaneously perform logical error suppression and to be robust DD sequences at the physical level, resistant to both control errors and crosstalk. Our results, using the $[[4,2,2]]$ code and IBM transmon qubits, demonstrate a beyond-breakeven fidelity of entangled logical qubits.
Our findings address a need along the path toward fault-tolerant quantum computation: keeping codes relatively small and nimble while still effectively handling logical errors.
Future research should aim to optimize LDD sequences tailored to specific codes and integrate QEC-LDD into quantum algorithms. Another interesting future direction is optimization of QEC-LDD for tunable-coupler transmon devices; we present preliminary results in Appendix E.
Acknowledgements.
This material is based upon work supported by, or in part by, the Intelligence Advanced Research Projects Activity (IARPA), under the Entangled Logical Qubits program through Cooperative Agreement Number W911NF23-2-0216, by the U.S. Army Research Laboratory and the U.S. Army Research Office under contract/grant number W911NF2310255, and by the Defense Advanced Research Projects Agency under Agreement HR00112230006. The views, opinions and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This research was conducted using IBM Quantum Systems provided through University of Southern California’s IBM Quantum Innovation Center. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A [[4,2,2]] code
@@ Acknowledgements.
Our particular choice of logical operators is motivated by the fact that $\overline{XX}$ and $\overline{YY}$ (i.e., the logical operators we use to implement the LDD sequences) have a natural staggering of their physical $X$ and $Y$ gates. In other words, in the implementation of the logical operators comprising LDD, nearest-neighbor qubits are always interleaved with an identity operation (e.g., $IXIX$ as opposed to $IXXI$). This is critical for nearest-neighbor crosstalk cancellation [65].

### v2

@@ IV.2 Logical error suppression and detection by NDD
For all our experimental results, “No DD” refers to $[[4,2,2]]$ encoding without any physical or logical DD. In all other experiments, we use physical XY4 to pad all the encoding and unencoding idle gaps, and combine them with various flavors of NDD. This choice allows us to clearly assess the improvements introduced by DD, compared to relying solely on the code’s error detection capabilities.
IV.2.1 Logical $Z$ errors
@@
We start by gradually increasing the time delay $\tau$ (up to 55 $\mu$s) between the encoding and unencoding without DD. This situation is relevant in the context of QEC experiments. For example, one could prepare an encoded qubit and then leave it to idle while other logical operations are applied to other encoded qubits [19].
We perform the $\overline{Z}_{i}$ error detection using the logical Bell states |$\ket{\Psi_{+}}$ and |$\ket{\Phi_{+}}$ as discussed in Section III, and display the results in Fig. 3. As can be seen in Fig. 3(a) and (c), without NDD the free evolution fidelity (denoted as No DD) decays rapidly and exhibits $ZZ$ crosstalk-induced oscillations. Figure 3(b) and (d) show that logical $Z$ errors accumulate over time.
@@
We next use two types of NDD sequences to suppress this effect: Normalizer XX (NXX) and Normalizer XY4 (NXY4). Physical-level schematics of these NDD sequences are shown in the boxed four-pulse sequences of Fig. 1(d) and (e), respectively (disregarding the tilde notation). Their logical-level counterparts are shown at the bottom of Fig. 1. We generate both NXX and NXY4 using the native logical operations of the $[[4,2,2]]$ code. For example, in the upper part of Fig. 1(d), reading the first (second) column from top to bottom yields $XIXI=\overline{XX}$ ($\overline{X^{\prime}X^{\prime}}=IXIX=S_{1}\overline{XX}$; $S_{1}=XXXX$), which is the $\bar{X}\bar{X}$ ($\bar{X}^{\prime}\bar{X}^{\prime}$) column at the bottom of Fig. 1(d). Reading the pulse sequences from left to right, the staggering (appearance of delays as indicated by the identity operations) is deliberately introduced to suppress crosstalk at the physical level [65]; see Appendix E for more details.
@@
The corresponding results are denoted as NXX (red) and NXY4 (dark green) in Fig. 3. NXX is a single-axis non-universal sequence, while NXY4 is “nearly” universal: it suppresses all logical errors except its own generating set $\{XIXI,IXIX,YIYI,IYIY\}$. At short times ($\lesssim 15\mu$s), these results are better than those without DD because some errors (including $ZZ$ crosstalk) are suppressed by both NXX and NXY4. However, at longer times, the benefit is lost and, moreover, small oscillations appear that indicate the presence of coherent errors  [89]. To overcome this, we create robust versions of NXY4 and NXX by ensuring that all physical qubits in the code undergo physical DD sequences robust to pulse errors. Specifically, we use the universally robust (UR) sequence family [62], and ensure that each physical qubit undergoes a UR4 sequence, i.e., $X\tilde{X}\tilde{X}X$ or $Y\tilde{Y}\tilde{Y}Y$, where a tilde denotes an $X$ or $Y$ rotation by $-\pi$ instead of $\pi$. These robust versions, which we call RNXX and RNXY4, are the full sequences shown in Fig. 1(d) and (e). The performance of RNXX (yellow) and RNXY4 (blue), as seen in Fig. 3, exhibits a significant improvement. Notably, the logical Bell state fidelities decay more slowly and without oscillations [panels (a) and (c)], and the logical $Z$ errors are strongly suppressed [panels (b) and (d)].
@@
In addition, we apply the physical staggered XY4 (SXY4) sequence. This corresponds to applying a single XY4 sequence to each physical qubit but in a staggered manner to reduce crosstalk [65]. This sequence, shown in Fig. 1(f), also corresponds to a non-universal purely logical-$Z$-error suppressing, NDD sequence for the $[[4,2,2]]$ code [bottom of panel (f)]. Interestingly, Fig. 3 shows that SXY4 performs on par with RNXX and RNXY4. This finding signals that the dominant logical errors are of $Z$-type. We confirm this in the following.
@@ IV.2.2 Logical $X$ and $Y$ errors
IV.2.3 NDD with postselection
Having implemented NDD, we can further improve the results by using the error detection capability of the $[[4,2,2]]$ code, which allows us to perform postselection. We do so by discarding any measurement outcome outside the logical basis (which would be the result of a physical error); i.e., we only keep measurement results corresponding to the bitstrings $\{0000,0110,1001,1111\}$. Doing so leads to our central experimental result, the demonstration of high-fidelity entangled logical qubits, shown in Fig. 5.
@@ IV.2.3 NDD with postselection
IV.2.4 Beyond breakeven and the state-of-the-art
Fig. 5 also includes results for physical (unencoded) Bell states. Here we show only the best Bell pair among all pairs we tested, both without DD and with staggered (crosstalk-robust [65]) XY4. The former (‘Best Physical No DD’) exhibits strong crosstalk-induced oscillations with an overall fidelity comparable to that of the mean logical encoded Bell pairs case without DD (No DD+PS). Adding SXY4 significantly improves the fidelity and outperforms even the mean fidelity of logical Bell pairs with NDD (RNXY4). This shows that NDD by itself is not better than working with physical qubits and a crosstalk-robust DD sequence. However, physical Bell pairs with SXY4 are significantly worse than RNXY4+PS, i.e., the case of NDD with postselection on the results of the $[[4,2,2]]$ code. This constitutes clear evidence of beyond-breakeven performance for our QEC-NDD strategy.
@@ IV.2.4 Beyond breakeven and the state-of-the-art
Figure 6: The percentage of data discarded per circuit as a function of the delay time between encoding and unencoding (dataset 2). While NDD suppresses logical errors, it also reduces certain physical errors, as evidenced by the lower data discard rate when using NDD (i.e., fewer physical errors are detected).
IV.3 Physical error suppression and detection by NDD
As explained in the discussion following 1 and in Section II.4, the NDD sequences suppress not only logical errors but physical errors as well. To see this, consider, e.g., the RNXX sequence [Fig. 1(d)]. As described in Section IV.2.1, reading the pulse sequence vertically, each time step of this sequence operates in the logical subspace as $\overline{XX}$. Simultaneously, reading the sequence horizontally over a complete round of NDD, all four physical qubits undergo the physical XX sequence (i.e., $X-\tau-X-\tau$), which suppresses the set of physical errors $\{Y_{i},Z_{i}\}_{i=1}^{4}$. Similarly, for the RNXY4 sequence [Fig. 1(e)], each physical qubit undergoes UR4. which robustly [62] suppresses the sets of physical errors $\{Y_{i},Z_{i}\}_{i=1,3}$ (due to the $X\tilde{X}\tilde{X}X$ sequence) and $\{X_{i},Z_{i}\}_{i=2,4}$ (due to the $Y\tilde{Y}\tilde{Y}Y$ sequence).
One way to gauge the impact of this suppression of physical errors is shown in Fig. 6, which displays the percentage of discarded data per circuit as a result of postselection over $55\mu$s (out of $4000$ shots), comparing No DD to RNXY4 for the two logical Bell states |$\ket{\Phi_{+}}$ and |$\ket{\Psi_{+}}$. The results are obtained through bootstrapping across the ten sets of four qubits used in these experiments. Significantly less data is discarded with RNXY4 than without DD. Since discarded data corresponds to the detection of physical errors, this means that NDD not only enhances the fidelity of logical Bell states by suppressing logical errors, but also reduces the occurrence of physical errors.
We can go further and use Algorithmic Error Tomography [54] to identify specific physical error types. For example, the $0100$ and $0010$ bitstrings correspond, respectively, to physical $Z_{2},Z_{4}$ and $Z_{1},Z_{3},$ errors in the |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ logical Bell state experiment. In Fig. 7 we show the corresponding relative bitstring counts, which are measures of detecting either of these errors. Additionally, we show the $0110$ and $0000$ relative bitstring counts, which correspond to the logical fidelity and the logical $Z$ error, respectively. Without DD (dashed), we observe that the logical fidelity is low, and both logical and physical errors increase and oscillate. With DD (solid), the RNXY4 sequence strongly suppresses both types of errors.
These results demonstrate that, as claimed, NDD sequences suppress both physical and logical errors.
Figure 7: The relative count of the $\{0000,0110,0100,0010\}$ bitstrings in the |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ experiments (dataset 2). The two even-weight bitstrings correspond to the logical fidelity ($0110$) and logical $Z$ error ($0000$), and the two odd-weight bitstrings correspond to the detection of physical $\{Z_{i}\}_{i=1}^{4}$ errors. Dashed (solid) lines correspond to no DD (RNXY4). The NDD sequence, in addition to suppressing the logical errors and improving the logical fidelity, also strongly suppresses the occurrence of these physical errors.
V Discussion
The operation of QEC codes is adversely affected by the occurrence of logical errors that the code cannot detect or correct. Here, we have shown how to combine QEC with dynamical decoupling implemented in terms of the normalizer elements of the code, resulting in a hybrid QEC-NDD strategy that is significantly more effective than either QEC or NDD alone. We designed our NDD sequences to simultaneously perform logical error suppression and to be robust DD sequences at the physical level, resistant to both control errors and crosstalk. As an added benefit, these sequences suppress many physical errors as well. Our results using the $[[4,2,2]]$ code and IBM transmon qubits, featured in Fig. 5, demonstrate a beyond-breakeven fidelity of entangled logical qubits. The fidelities we report are the highest to date for entangled logical qubits using superconducting qubits.
Our findings address a need along the path toward fault-tolerant quantum computation: keeping codes relatively small and nimble while still effectively handling logical errors. Future research should aim to optimize NDD sequences tailored to specific codes and integrate QEC-NDD into quantum algorithms. Another interesting future direction is the optimization of QEC-NDD for tunable-coupler transmon devices; we present preliminary results in Appendix F.
Acknowledgements.
This material is based upon work supported by, or in part by, the Intelligence Advanced Research Projects Activity (IARPA), under the Entangled Logical Qubits program through Cooperative Agreement Number W911NF23-2-0216, by the U.S. Army Research Laboratory and the U.S. Army Research Office under contract/grant number W911NF2310255, and by the Defense Advanced Research Projects Agency under Agreement HR00112230006. The views, opinions and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This research was conducted using IBM Quantum Systems provided through the University of Southern California’s IBM Quantum Innovation Center. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A Quantum codes and logical errors
A general $[[n,k,d]]$ code encodes $k$ logical qubits into $n$ physical qubits with distance $d$ [92]. The weight $w$ of an error is the number of physical qubits it affects simultaneously. An $[[n,k,d]]$ stabilizer code corrects every error whose weight satisfies $w\leq t=\bigl{\lfloor}\tfrac{d-1}{2}\bigr{\rfloor}$ and detects every error whose weight $w\leq d-1$. A $d=2$ code is a pure quantum error detection (QED) code, and a distance-$d$ QEC code can always be used as a QED code for errors of weight $<d$. The code distance is the minimum number of physical qubits that must experience an error to cause an undetectable logical error, i.e., an error forming a logical operation inside the code space. These logical errors can either be inherently present or result from the accumulation of lower-weight errors over time. For example, in the context of superconducting qubits with fixed-frequency couplers, a prevalent challenge is $ZZ$ crosstalk [93, 94], which inherently introduces weight-two errors that can present as logical errors for distance-$2$ codes.
@@ Appendix A Quantum codes and logical errors
The normalizer (or centralizer) of $\mathcal{S}$ in $\mathcal{P}_{n}$ is the group $\mathcal{N}(\mathcal{S})$ of Pauli operators commuting with all elements of $\mathcal{S}$ [^2] The normalizer $\mathcal{N}(\mathcal{S})=\mathcal{N}_{\mathcal{G}}(\mathcal{S})$ of $\mathcal{S}\subseteq\mathcal{G}$ is defined as $\{g\in\mathcal{G}\text{ s.t. }g\mathcal{S}g^{-1}=\mathcal{S}\}$, and the centralizer $\mathcal{Z}(\mathcal{S})=\mathcal{Z}_{\mathcal{G}}(\mathcal{S})$ of $\mathcal{S}\subseteq\mathcal{G}$ is defined as $\{g\in\mathcal{G}\text{ s.t. }\forall s\in\mathcal{S},gs=sg\}$. These two notions coincide in the case of stabilizer groups. The normalizer is also uniquely defined. However, the group of canonical logical operators is not unique: one can choose arbitrary stabilizers and multiply the canonical generators $\bar{X}_{j},\bar{Z}_{j}$ by those to obtain other choices of the group $\mathcal{L}$.
Another important set is $\mathcal{D}=\mathcal{P}_{n}\setminus\mathcal{N}(\mathcal{S})$, the set of detectable errors. Since every pair of elements of $\mathcal{P}_{n}$ either commute or anticommute, and $\mathcal{N}(\mathcal{S})$ contains all elements of $\mathcal{P}_{n}$ that commute with $\mathcal{S}$, every element of $\mathcal{D}$ must anticommute with at least one element of $\mathcal{S}$.
Appendix B [[4,2,2]] code
@@
Our particular choice of logical operators is motivated by the fact that $\overline{XX}$ and $\overline{YY}$ (i.e., the logical operators we use to implement the LDD or NDD sequences) have a natural staggering of their physical $X$ and $Y$ gates. In other words, in the implementation of the logical operators comprising NDD, nearest-neighbor qubits are always interleaved with an identity operation (e.g., $IXIX$ as opposed to $IXXI$). This is critical for nearest-neighbor crosstalk cancellation [65].